The perimeter of a rectangular volleyball court is 180 feet.The court's width,w,is half of its length,l.Determine the deminsions,in feet,of the volleyball court.
Answer by math_helper(2461)

2L + 2W = 180
and
L=2W
so
2(2W) + 2W = 180
4W + 2W = 180
6W = 180
+W+=+30ft+ and +L+=+60ft+